OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. (9, -9);8x-3y=5

OpenStudy (anonymous):

I would first rewrite the equation given in slope-intercept form (y=mx+b), and the slope of your answer will be the same as the slope in the given equation. Then, you can plug in the x and y values into your new equation and then solve for b. This will give you the b value.

OpenStudy (anonymous):

my slope came out to be 8/3 which makes the equation y=8/3x + 5. the problem only asks for the equation so would this be right or did i mess up somewhere along the way

OpenStudy (anonymous):

I don't know if the figures are right, you might want to double check your work if you're unsure. But, that would be a perfectly acceptable format. if it only wants the equation.

OpenStudy (anonymous):

Kizzfan911 made a mistake. you get the equation y = (8/3)x - (5/3). Now we know the slope so plug in the point given into y = (8/3)x + b and solve for b giving you the answer.

OpenStudy (anonymous):

i came out with -33 so if smstevns is correct in his steps that would make the answer y = (8/3)x + 33 right?

OpenStudy (anonymous):

sorry y = (8/3)x - 33

OpenStudy (anonymous):

I just worked out the problem, and that's the answer I got. Looks good.

OpenStudy (anonymous):

thank you kuromeru

OpenStudy (anonymous):

No problem, happy to help.